OpenStudy (anonymous):

1. Euros are currency used in several European countries. On one day in October 2005, you could exchange $3 for about 2.5 euros. How many dollars would you have needed to get 8 Euros? 2. Hooke’s law states that the distance a spring is stretched is directly proportional to the force applied. If 20 pounds of force stretches a spring 4 inches, how much will the spring stretch if 80 pounds of force is applied? 3. A 3-ounce serving of tuna fish provides 24 grams of protein. How many grams of protein are in a 10-ounce serving of tuna fish?

OpenStudy (bibby):

\(\dfrac{$3}{2.5Euro}=\dfrac{$x}{8euro}\)

OpenStudy (bibby):

\(\dfrac{20lb}{4in}=\dfrac{80lb}{xin}\) \(\dfrac{3oz}{24g}=\dfrac{10oz}{xg}\) do you need help solving these?

OpenStudy (anonymous):

no, thank you

OpenStudy (bibby):

alrighty, feel free to post answers if you want them confirmed

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yA i need help after all ;///

OpenStudy (bibby):

:\\\\

OpenStudy (bibby):

:V :|

OpenStudy (bibby):

although technically, "cross multiplying" isn't a valid operation, it's the commonly used method to teach people how to solve proportions.